### v1

# Making Machines Sound Sarcastic: LLM-Enhanced and Retrieval-Guided Sarcastic Speech Synthesis

Abstract
Sarcasm is a subtle form of non-literal language that poses significant challenges for speech synthesis due to its reliance on nuanced semantic, contextual, and prosodic cues. While existing speech synthesis research has focused primarily on broad emotional categories, sarcasm remains largely unexplored. In this paper, we propose a Large Language Model (LLM)-enhanced Retrieval-Augmented framework for sarcasm-aware speech synthesis. Our approach combines (1) semantic embeddings from a LoRA-fine-tuned LLaMA 3, which capture pragmatic incongruity and discourse-level cues of sarcasm, and (2) prosodic exemplars retrieved via a Retrieval Augmented Generation (RAG) module, which provide expressive reference patterns of sarcastic delivery. Integrated within a VITS backbone, this dual conditioning enables more natural and contextually appropriate sarcastic speech. Experiments demonstrate that our method outperforms baselines in both objective measures and subjective evaluations, yielding improvements in speech naturalness, sarcastic expressivity, and downstream sarcasm detection.
Index Terms—  Sarcastic speech synthesis, sarcasm detection, large language models, retrieval augmented generation
1 Introduction
Sarcasm is a common yet challenging form of non-literal language, often marked by prosodic patterns such as exaggerated intonation and prolonged syllables [1, 2]. Its proper interpretation relies on a nuanced interplay of semantic, contextual, and prosodic cues, making it difficult to capture computationally. While sarcasm is widely used in everyday communication, it has received limited attention in the field of speech synthesis. Most existing work on expressive text-to-speech (TTS) has concentrated on broad emotional categories such as happiness, anger, or sadness [3, 4, 5], whereas more subtle and pragmatically grounded attitudes such as sarcasm remain largely unexplored. While prior work has attempted to analyze acoustic correlates of sarcastic speech, only a limited number of studies have explored directly manipulating prosodic features such as pitch, pace, and loudness to generate sarcastic-sounding speech [6]. In addition, current TTS systems, while successful at producing intelligible and emotionally expressive speech, often lack the nuanced expressiveness needed to convey sarcastic intent [5, 7, 8].
Sarcastic speech synthesis holds considerable potential for improving human-computer interaction in applications such as conversational agents and entertainment systems [9]. However, the task presents two major challenges. First, high-quality, annotated sarcastic speech corpora remain scarce, limiting the applicability of purely data-driven approaches. Existing corpora for sarcasm, such as MUStARD [10] or its extensions [11], are primarily developed for detection rather than synthesis. Second, sarcasm is inherently more subtle and context-dependent than conventional emotions such as anger or joy, making it difficult to capture with handcrafted acoustic features or simple prosodic controls. In contrast, research on sarcasm detection has progressed significantly. Recent work has shown that multi-modal approaches, incorporating textual, prosodic and visual information, can significantly improve sarcasm recognition [10, 11, 12, 13, 14]. These advances highlight the feasibility of modeling sarcasm computationally. Yet, most efforts have remained on the recognition side, leaving the generation of sarcastic speech largely unaddressed. Moving from detection to synthesis is not only a natural next step but also a crucial one, as it enables interactive systems to actively deploy pragmatic phenomena rather than passively identify them.
To address these limitations, recent advances in natural language processing offer promising directions. LLMs have demonstrated the ability to capture high-level semantic and pragmatic features from text. In the TTS domain, integrating LLM-derived embeddings has been shown to enable more nuanced prosody control and improve emotional expressiveness and contextual appropriateness [15, 16, 17]. Such capabilities are particularly relevant for sarcasm, which often relies on pragmatic incongruity between literal meaning and intended attitude. Leveraging LLM-based semantic representations could therefore provide a principled way of conditioning TTS models on sarcasm-relevant cues that handcrafted features fail to capture.
Complementary to this, RAG allows models to condition generation on external knowledge or examples retrieved from large databases [18, 19]. While RAG has primarily been explored in text-based tasks such as question answering and dialogue [20, 21], its potential for speech synthesis is evidenced by related work on style transfer and expressive TTS, where reference utterances are used to guide prosody [18]. Existing approaches, however, typically rely on manually selected or speaker-specific reference audio, limiting scalability and contextual alignment. In the case of sarcasm, where annotated corpora are scarce, automatic retrieval of semantically similar sarcastic speech could provide prosodic exemplars that both enrich training and guide generation. This suggests a paradigm in which LLMs supply semantic-pragmatic cues while RAG retrieves expressive references, together enabling a systematic modeling of subtle pragmatic phenomena such as sarcasm in speech synthesis.
Fig. 1: Overview of the sarcasm-aware speech synthesis framework. Semantic cues from a LoRA fine-tuned LLaMA 3 and prosodic exemplars retrieved via RAG are fused within a VITS-based model to synthesize sarcastic speech.
This study introduces a unified sarcasm-aware speech synthesis framework that enriches the TTS system with LLM-extracted semantic cues indicative of sarcasm and prosodically aligned audio references for expressive modeling. Specifically, we first use an LLM to extract high-level semantic representations capturing features relevant to sarcastic intent. In parallel, a retrieval mechanism based on RAG identifies semantically similar and contextually aligned prosodic examples from a curated speech database, providing prosodic exemplars that guide the TTS model toward expressive and contextually appropriate output. Both objective metrics and subjective evaluations demonstrate that the proposed method outperforms baseline models in naturalness, expressivity, and accurate sarcasm expression.
2 Methods
To achieve high-quality sarcastic speech synthesis, we propose a Retrieval-Augmented LLM-enhanced TTS framework. An overview of the architecture is illustrated in Figure 1.
2.1 Sarcasm-Aware Semantic Encoding via LoRA
Low-Rank Adaptation (LoRA) has recently been widely adopted for adapting LLMs across diverse domains such as emotion recognition, sentiment analysis, and affective detection, due to its parameter efficiency and effectiveness [22]. Motivated by these advances, we leverage LoRA to finetune the LLaMA 3-8B model for sarcasm-aware semantic encoding. Fine-tuning is performed on a curated sarcasm-labeled dataset, enabling the model to capture signals of sarcastic intent, including pragmatic incongruity and discourse-level cues.
Formally, given an input text $x$, the Sarcasm-LoRA encoder produces a semantic embedding:
$\mathbf{E}_{s}=f_{\text{LoRA}}(x)\in\mathbb{R}^{T_{t}\times d_{t}},$ (1)
where $d_{t}$ denotes the dimensionality of the semantic embedding. These embeddings capture both contextual and pragmatic aspects of the input, providing a sarcasm-aware representation for further processing.
2.2 RAG for Prosody Conditioning
To enhance prosodic expressiveness, we integrate a RAG module into the synthesis pipeline. The key idea is to leverage semantically aligned sarcastic utterances as prosodic exemplars, enriching the generated speech with realistic intonation patterns. Concretely, we first construct an index $\mathcal{D}=\{(u_{i},\mathbf{a}_{i})\}_{i=1}^{N}$ of sarcastic utterances $u_{i}$ curated from MUStARD++, where $\mathbf{a}_{i}$ denotes their pre-computed semantic representations. For a given input text, the semantic embedding $\mathbf{E}_{s}$ produced by the Sarcasm-LoRA encoder is used as a retrieval query to fetch the top-$K$ semantically and pragmatically relevant sarcastic utterances from the indexed database.
We first compute the cosine similarity between $\mathbf{E}_{s}$ and each database entry:
$\text{sim}(\mathbf{E}_{s},\mathbf{a}_{i})=\frac{\mathbf{E}_{s}\cdot\mathbf{a}_{i}}{\|\mathbf{E}_{s}\|\|\mathbf{a}_{i}\|}.$ (2)
We then retrieve the top-$K$ most relevant utterances. Each retrieved utterance $u_{k}\in\mathcal{U}_{\text{top-}K}$ is encoded by WavLM [23] to obtain a prosody embedding:
$\mathbf{E}_{w_{k}}=\text{Pool}\!\left(f_{\text{WavLM}}(u_{k})\right)\in\mathbb{R}^{d_{w}},$ (3)
where $\text{Pool}(\cdot)$ aggregates the frame-level features into a fixed-length vector. These exemplars provide style references that reflect the characteristic intonation and emphasis patterns of sarcastic speech.
Compared to conventional style transfer methods that rely on a single manually chosen reference utterance, the proposed retrieval mechanism offers two key advantages. First, it scales more naturally to diverse input contexts, since the retrieval process automatically identifies semantically aligned exemplars. Second, by conditioning on multiple retrieved samples, the model can capture a richer variety of sarcastic prosodic patterns, leading to more robust and contextually appropriate expressive synthesis.
2.3 LLM-Enhanced and Retrieval-Guided Sarcastic TTS
The system builds upon the VITS architecture [24], enriched with semantic features extracted by a fine-tuned LLaMA 3 and guided by prosodic exemplars retrieved via a RAG module.
Let $\mathbf{E}_{p}\in\mathbb{R}^{T_{p}\times d_{p}}$ denote the phoneme embeddings and $\mathbf{E}_{s}\in\mathbb{R}^{T_{t}\times d_{t}}$ the sarcasm-aware semantic embeddings produced by the fine-tuned LLaMA 3. In the cross-attention module, we treat $\mathbf{E}_{p}$ as the query and $\mathbf{E}_{s}$ as key and value:
$\mathbf{Q}=\mathbf{E}_{p}\mathbf{W}_{q},\quad\mathbf{K}=\mathbf{E}_{s}\mathbf{W}_{k},\quad\mathbf{V}=\mathbf{E}_{s}\mathbf{W}_{v},$ (4)
$\mathbf{H}=\text{Softmax}\Big(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d_{k}}}\Big)\mathbf{V},$ (5)
where $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$ are learned projection matrices, and $d_{k}$ is the dimensionality of the key.
Finally, the prosody exemplars $\mathcal{E}_{w}=\{\mathbf{E}_{w_{1}},\dots,\mathbf{E}_{w_{K}}\}$ extracted by WavLM are linearly projected and added to the cross-attention output to modulate the decoder:
$\mathbf{Z}=\mathbf{H}+\sum_{k=1}^{K}\mathbf{W}_{w}\mathbf{E}_{w_{k}},$ (6)
where $\mathbf{W}_{w}$ is a learned linear projection mapping the prosody embeddings to the decoder dimension. This yields hidden states $\mathbf{Z}$ that are conditioned on phonemes, semantic content, and retrieved prosodic cues, enabling the synthesis of sarcastic speech with expressive intonation.
3 Experiments
To evaluate the effectiveness of our proposed approach to sarcastic speech synthesis, we conducted a series of experiments involving baseline, LLM-enhanced, RAG-enhanced, and combined models. These experiments assess the impact of semantic guidance from language models and reference-based prosodic conditioning on the quality and sarcasm-awareness of synthesized speech [^1] Speech samples are available at https://abel1802.github.io/RAG-LLM-SarcasticTTS/.
3.1 Data
For the initial pre-training phase of VITS, we use the HiFi-TTS corpus [25], a high-fidelity audiobook-derived dataset with diverse speakers and wideband, clean recordings.
To build the sarcastic speech database for retrieval purposes, we use the MUStARD++ dataset [11]. MUStARD++ is a multimodal sarcasm detection corpus compiled from popular sitcoms such as Friends and The Big Bang Theory. It contains 1,202 audiovisual utterances, equally divided into 601 sarcastic and 601 non-sarcastic samples. Following standard practice, we partition the corpus into training, validation, and test sets with an 8:1:1 ratio. Specifically, 10% of the data is held out for evaluation, while the remaining 80% of sarcastic samples are used to construct the retrieval-based sarcastic speech database. This setup ensures relatively sufficient coverage of sarcastic prosody in the retrieval dataset.
To adapt LLMs as sarcasm-aware semantic encoders for speech synthesis, we fine-tune them on the News Headlines Sarcasm dataset [26]. This dataset contains over 28,000 headlines from sources such as The Onion, each annotated with a binary sarcasm label. The class distribution is roughly balanced, making it suitable for learning discourse-level markers of sarcastic intent. The LLM serves as a feature extractor that provides high-level semantic embeddings conditioned on sarcastic intent. To validate the effectiveness of this adaptation, we also evaluate the fine-tuned model on a sarcasm detection task on the whole MUStARD++ dataset [11], in order to test that the learned embeddings capture sarcasm-relevant cues.
3.2 Experimental Setup
For training the baseline VITS model, we follow the standard configuration described in [24]. The model is pre-trained on the HiFi-TTS corpus [25] using the open-source Amphion toolkit [^2] https://github.com/open-mmlab/Amphion.
To adapt LLMs for sarcasm-aware synthesis, the LoRA fine-tuning of LLaMA 3 is performed with an expansion factor of 8 and a learning rate of 1e41e-4, updating only the low-rank adapters in the attention layers while freezing the backbone weights [^3] https://github.com/hiyouga/LLaMA-Factory.
3.3 Compared Methods
We compare the proposed framework against a series of baseline and enhanced variants to systematically evaluate the contribution of different modules:
  • VITS (Baseline): A standard variational text-to-speech model serving as the baseline, without explicit semantic or stylistic guidance.
  • VITS + BERT (w/ BERT): Extends VITS with semantic embeddings from a pretrained BERT model [27], providing additional semantic features during synthesis.
  • VITS + LLaMA 3 (w/ LLaMA 3): Uses textual embeddings generated by LLaMA 3 [28], enabling richer contextual and pragmatic guidance.
  • VITS + LLaMA 3-LoRA (w/ LoRA): Incorporates a LoRA-fine-tuned version of LLaMA 3 trained on sarcasm-labeled data, enhancing the ability to capture sarcasm-relevant semantics.
  • VITS + RAG (w/ RAG): Employs RAG to retrieve sarcastic reference speech samples that are semantically similar to the input text. These references are used to provide prosodic guidance during synthesis.
  • VITS + LLaMA 3-LoRA + RAG (Proposed): Our proposed unified framework, combining sarcasm-aware semantic embeddings from fine-tuned LLaMA 3 with prosodic exemplars retrieved through RAG. This model is designed to maximize both semantic relevance and expressive alignment with sarcastic intent.
4 Results and Discussion
4.1 Sarcasm Detection Results
We evaluate sarcasm detection performance from text-only inputs on the MUStARD++ dataset, comparing the original sarcasm detection systems used in MUStARD++ [11], BERT, LLaMA 3, and LLaMA 3-LoRA. Models are evaluated with Precision (P), Recall (R), and weighted F1-score (F1), with results summarized in Table 1.
Table 1: Performance of different models on sarcasm detection (text-only input).
Method P (%) R (%) F1 (%)
MUStARD++ (text) 67.9 67.7 67.7
BERT 66.8 66.9 66.8
LLaMA 3 65.7 65.4 65.5
LLaMA 3-LoRA 72.4 72.7 72.5
Table 2: Evaluation results of TTS systems. MCD is in dB; pitch and energy are averaged with standard deviation; Sarcasm detection includes precision (P), recall (R), and weighted F1-score (F1); MOS and SMOS denote naturalness and sarcasm MOS scores, respectively.
Objective Sarcasm Detection Subjective
Method MCD $\downarrow$ Pitch $\downarrow$ Energy $\downarrow$ P (%) $\uparrow$ R (%) $\uparrow$ F1 (%) $\uparrow$ MOS $\uparrow$ SMOS $\uparrow$
GT 62.8 62.3 62.3 $3.8\pm 0.1$ $4.5\pm 0.2$
Baseline $9.8\pm 3.0$ $261.9\pm 148.8$ $4.4\pm 1.4$ 60.3 60.5 59.9 $2.6\pm 0.1$ $3.2\pm 0.2$
w/ BERT $\mathbf{9.6\pm 2.8}$ $265.6\pm 147.1$ $4.5\pm 1.4$ 60.5 60.6 60.5 $2.5\pm 0.1$ $3.1\pm 0.2$
w/ LLaMA 3 $10.4\pm 3.1$ $261.6\pm 135.1$ $4.4\pm 1.0$ 59.6 59.7 59.0 $2.0\pm 0.1$ $2.6\pm 0.2$
w/ LoRA $10.1\pm 3.1$ $282.6\pm 149.6$ $4.4\pm 1.4$ 60.6 60.8 60.6 $2.6\pm 0.1$ $3.8\pm 0.2$
w/ RAG $10.0\pm 3.0$ $261.0\pm 148.0$ $4.4\pm 1.4$ 61.5 61.7 61.6 $2.6\pm 0.1$ $3.7\pm 0.2$
Proposed $9.8\pm 2.8$ $\mathbf{259.6\pm 152.3}$ $\mathbf{4.4\pm 1.4}$ 62.7 62.9 62.5 $\mathbf{2.7\pm 0.1}$ $\mathbf{3.8\pm 0.2}$
As shown in Table 1, both the MUStARD++ baseline and BERT achieved moderate performance with F1-scores around 67%. LLaMA 3 performed slightly worse (65.5%), suggesting that general-purpose pretraining alone is insufficient to capture the nuanced cues of sarcasm. In contrast, parameter-efficient fine-tuning with LoRA yielded a clear improvement, raising the F1-score to 72.5%. This demonstrates the effectiveness of the proposed sarcasm-aware semantic encoding via LoRA (Section 2.1), and motivates its integration into the sarcasm-aware TTS framework to further examine whether such improvements in semantic encoding can also lead to more expressive synthesized speech.
4.2 Sarcastic Speech Synthesis Results
Table 2 summarizes the results of different TTS systems across three perspectives: low-level acoustic metrics (MCD, pitch, energy), downstream sarcasm detection performance, and subjective naturalness ratings (Natural MOS and Sarcasm MOS). To assess downstream sarcasm detection performance, we adopt the detection architecture used in MUStARD++ with collaborative gating [11], using only the speech modality. Features extracted from synthesized audio are compared against target sarcasm labels, where higher agreement indicates stronger sarcasm expressiveness.
4.2.1 Objective Evaluation
Overall, the proposed VITS + LLaMA 3-LoRA + RAG system achieves the strongest performance across both dimensions. It obtains the lowest distortion ($9.83$ MCD) and the most stable prosodic statistics, while also yielding the highest sarcasm detection weighted F1-score (62.5%). In comparison, VITS + LLaMA 3 suffers from degraded quality (higher MCD), suggesting that raw embeddings from an untuned LLM do not align well with the synthesis task. LoRA fine-tuning mitigates this issue, and retrieval-augmented guidance further enhances both acoustic stability and sarcasm classification accuracy.
Across all systems, acoustic metrics such as MCD, pitch, and energy remain relatively stable, indicating that the improvements stem less from low-level signal fidelity and more from high-level semantic and prosodic conditioning.
4.2.2 Subjective Evaluation
In addition to objective metrics, we conducted a subjective evaluation to assess the perceptual quality and sarcastic expressiveness of the synthesized speech (Table 2). We recruited 30 listeners from diverse backgrounds who were asked to rate randomly shuffled samples from each system on two dimensions: (1) naturalness, using the standard Mean Opinion Score (MOS) on a 5-point Likert scale, and (2) sarcasm expressiveness (SMOS), reflecting how well the output conveyed sarcastic intent.
Pairwise comparison across systems reveals several important trends. First, the baseline VITS system received moderate MOS ($2.6\pm 0.1$) and SMOS ($3.2\pm 0.2$), indicating intelligible but relatively flat speech with limited sarcastic cues. When replacing the text encoder with BERT, scores remained similar ($2.5\pm 0.1$ MOS, $3.1\pm 0.2$ SMOS), suggesting that in our specific setting, BERT embeddings do not provide meaningful advantages for sarcasm conditioning.
In addition, directly integrating LLaMA 3 degraded performance ($2.0\pm 0.1$ MOS, $2.6\pm 0.2$ SMOS). Listeners frequently reported that these samples sounded less natural and had flat intonation patterns. This indicates that raw LLM embeddings are poorly aligned with intended expressive speech.
LoRA fine-tuning substantially narrowed this gap, raising naturalness to $2.6\pm 0.1$ and sarcasm expressiveness to $3.8\pm 0.2$. Compared to the baseline, listeners consistently noted clearer prosodic cues (e.g., exaggerated intonation) that aligned with sarcastic intent. This suggests that LoRA adaptation enabled LLaMA 3 embeddings to contribute pragmatically relevant sarcastic cues. Incorporating RAG provided a noticeable improvement ($2.6\pm 0.1$ MOS, $3.7\pm 0.2$ SMOS). Pairwise comparisons highlighted that RAG samples exhibited prosodic patterns more consistent with real sarcastic exemplars, making sarcastic intent more immediately recognizable.
Finally, the proposed system achieved the strongest results with $2.7\pm 0.1$ MOS and $3.8\pm 0.2$ SMOS. Taken together, the results suggest three main insights: (i) raw LLM embeddings are insufficient and may harm synthesis quality; (ii) parameter-efficient adaptation with LoRA can effectively inject pragmatic knowledge into the speech synthesis pipeline; and (iii) retrieval-augmented conditioning further refines prosodic expressiveness by grounding synthesis in real sarcastic exemplars. Together, these mechanisms enable the proposed system to generate speech that is both natural and perceptually recognizable as sarcastic.
5 Conclusion
We presented a unified framework for sarcasm-aware speech synthesis that leverages semantic conditioning from finetuned LLMs and prosodic guidance from retrieval-based exemplars. By fine-tuning LLaMA 3, our system captures semantic-level markers of sarcastic intent, while RAG provides prosodic references that enhance expressive alignment. Our results demonstrate the feasibility of modeling subtle pragmatic phenomena with TTS by combining parameter-efficient LLM adaptation and retrieval augmentation. In future work, we aim to extend this approach to other pragmatic styles such as humor, and explore cross-lingual sarcasm synthesis in diverse cultural contexts. We anticipate that this method could also serve as an effective data augmentation strategy for sarcasm detection tasks.